OpenStudy (anonymous):

Divide the complex numbers (3-2i)/(4+5i)

OpenStudy (anonymous):

is it 3/4?

OpenStudy (anonymous):

the choices are 1+25i/-9

OpenStudy (anonymous):

6+4i/18

OpenStudy (anonymous):

2-23i/41

OpenStudy (anonymous):

5-12i/41

OpenStudy (anonymous):

\[\frac{ 3-2i }{ 4+5i }*\frac{ 4-5i }{4-5i }\\\\ \frac{ ( 3-2i ) ( 4-5i ) }{ 4-5i }\\\\ \frac{2-23i }{ 41 } \]

OpenStudy (anonymous):

\[\Huge\infty \]

OpenStudy (anonymous):

Thanks ! :)